the mass of an aeroplane is 48000nkg.It starts from rest and develops a speed of 42 m/s just before takeoff.If the length of the runway is 2km,Calculate the power developed

Grade:9

1 Answers

Aarti Gupta
9 years ago
Dear student you have given the mass of aeroplane 48000nkg i think it should be 48000kg.
Given -
Mass of aeroplane = 48000Kg
Initial velocity ,u = 0
Final velocity , v = 42m/s
Length of runway i.e distance = 2Km = 2000m
Power,P = ?
Solution-
We know that, P = work done /time taken = W/t
Now,W = F * S ( force *distance)
And F = m * a (mass * acceleration)
To calculate a we use the formula, v = u + a*t
t,time taken = distance / speed.... ( as speed = distance/time)
t = 2000/42 = 47.62sec
v = 0 + a*t
a = v/t = 42/47.62 = 0.882m/s2
thus, F = 48000 * 0.882
and W = 48000 * 0.882 * 2000
W = 8.467 * 107J
P = 8.47 * 107 / 47.62
P = 1.778 * 106 watt or, 17.78 * 105watt